If all P-atom are removed from 490g H3PO4 and 2g molecules of ca(PO4)2 then find the moles and mass of p-atoms removed?

Grade:11

1 Answers

Chhavindra Tripathi
5 years ago
Moles of H3PO4 = 490/98=5
Atoms of P in 5 mole H3PO4= 5 mole atoms of P
And 2g molecules means 2 moles of Ca(PO4)2
It will contain 2*2 = 4 mole atoms of P
Total moles of P removed = 4+5= 9 mole P atoms
Mass removed = 9*31= 279 gm